### v2

# Advances in Kidney Biopsy Lesion Assessment through Dense Instance Segmentation

Abstract
Renal biopsies are the gold standard for diagnosis of kidney diseases. Lesion scores made by renal pathologists are semi-quantitative and exhibit high inter-observer variability. Automating lesion classification within segmented anatomical structures can provide decision support in quantification analysis and reduce the inter-observer variability. Nevertheless, classifying lesions in regions-of-interest (ROIs) is clinically challenging due to (a) a large amount of densely packed anatomical objects (up to 1000), (b) class imbalance across different compartments (at least 3), (c) significant variation in object scales (i.e. sizes and shapes), and (d) the presence of multi-label lesions per anatomical structure. Existing models lack the capacity to address these complexities efficiently and generically. This paper presents a generalized technical solution for large-scale, multi-source datasets with diverse lesions. Our approach utilizes two sub-networks: dense instance segmentation and lesion classification. We introduce DiffRegFormer, an end-to-end dense instance segmentation model designed for multi-class, multi-scale objects within ROIs. Combining diffusion models, transformers, and RCNNs, DiffRegFormer efficiently recognizes over 500 objects across three anatomical classes (glomeruli, tubuli, arteries) within ROIs on a single NVIDIA GeForce RTX 3090 GPU. On a dataset of 303 ROIs (from 148 Jones’ silver-stained renal WSIs), it outperforms state of art models, achieving AP of 52.1% (detection) and 46.8% (segmentation). Our lesion classification sub-network achieves 89.2% precision and 64.6% recall on 21889 object patches (from the 303 ROIs). Importantly, the model demonstrates direct domain transfer to PAS-stained WSIs without fine-tuning.
keywords:
MSC:
[…]
journal: Medical Image Analysis
1 Introduction
Expert pathologists’ evaluation of kidney biopsies remains the gold standard for diagnosing and staging renal diseases ([3]). Although biopsies digitalized into Whole Slide Images (WSIs) [^2] https://www.mbfbioscience.com/whole-slide-imaging-analysis/ have facilitated obtaining visual morphological assessment of different anatomical structures for disease categorization, high-quality diagnostic assessments heavily depend on the correct lesion quantification manually annotated by pathologists accross structures within a biopsy. Figure.1 shows an example of annotated region-of-interest (ROI) [^3] To prevent any confusion with ’RoI’ in RCNNs, we will hereafter use the term solely in the context of renal biopsies. Within RCNN discussions, we will adopt ’candidate regions’ or ’regional proposals’ to maintain clear distinctions. within a biopsy that contains hundreds of densely packed tissue objects. The annotation would cost a skilled expert around 2-4 hours for a complete biopsy. Due to the complexity and time-consuming nature of this task, there is a strong need for automated structure annotation and lesion classification tools to facilitate further quantification, offload annotation time, and reduce intra- / inter-observer variability ([1, 44]).
Deep learning based instance segmentation algorithms have demonstrated significant capababilities on biomedical datasets([47, 2, 28, 40, 15, 48]), but still face challenges to develop a generic framework in the domain of renal pathology due to 4 difficulties(refer to Figure.1): (1) densely packed structures per ROI (up to 1000); (2) vastly varying scales in objects (e.g. arteries can be up to 100 times larger than that of tubuli); (3) class unbalance (e.g. the tubulointerstitial area occupies more than 70% on average in healthy and diseased renal parenchyma ([2])); (4) ready for extensibility (i.e. cost-effectively adapt to new lesion types from expanding datasets in clinical scenarios). Simultaneously addressing these gaps, therefore, requires a universal framework with efficiency and flexibility.
Prior studies have shown significant capababilities in lesion classification through dense instance segmentation, but their paradigms lack scalability and adaptability to datasets with potential changes in lesion compositions. Some approaches ([23, 2, 40, 34, 15]) adopted a two-step process with semantic segmentation followed by dataset-specific post-processing to achieve final instance masks, which is unscalable to large-scaled datasets with dense objects in various shapes. Besides, each lesion is defined as one semantic class, leading to inflexible to potential changes in lesion combinations. Specifically, adding or removing lesions requires complete model redesign and retraining. Lastly, increasing lesion types must add more segmentation maps for prediction, which is significantly resource-intensive.
Detection-based models with regional convolution neural networks (RCNN) ([27, 28]) can efficiently detect dense anatomical structures and seperately design lesion classification head for each class. They can, therefore, adapt to lesion changes by replacing the corresponding sub-modules and maximally reuse the others. However, these variants are unscalable to multi-class objects at various scales and shapes, due to the reliance on pre-defined bounding box anchors, limiting their application to process single class with less variation in scales and shapes, like glomeruli.
Recently, transformers have been prevalent anchor-free approach to address multi-class objects at a variety of scales and shapes. Transformer-based instance segmentation utilizes attention mechanisms and learns latent representative embeddings (queries) from global contextual features to process vastly varying objects. However, existing models ([5, 12, 41]) are resource intensive for dense structures in large-scale datasets due to two drawbakcs: (1) large number of global queries from one embedding per object; (2) low instance map occupancy from one object per instance map (depicted in Figure.2. (c)). Hence, they are limited to process classes with sparse objects, like gomeruli.
Fig. 1: An illustration of a manually annotated ROI of kidney biopsy for glomeruli, tubuli, and arteries. There are 3 difficulties in clinical biopsies: (1) a large number of objects closely touch each other; (2) instances have a large variety in sizes and shapes; (3) the distribution of class is heavily biased.
In response to these challenges, We propose a generic and extensible system for dense instance segmentation and lesion classification on large-scale datasets with potential changes in lesion combinations. Our design has two key components: (a) A novel dense instance segmentation model recognises basic anatomical structures (glomeruli, tubuli, arteries); (b) a lesion classifier with a set of independent heads predicts lesions for each class. In a big picture, it is crucial to seperately modularize lesion classfication and dense instance segmentation. The segmentation of basic anatomical structures is an universal foundation for all visual assessment systems. But the lesion classification is task-driven down-stream application, which might vary with the interest of users. Therefore, the robustness of dense instance segmentaiton from local changes in heads of lesion classification is beneficial for the extensibility of our system to expanding datasets in clinical scenarios.
More specifically, for dense instance segmentation, we propose a novel end-to-end approach, named DiffRegFormer, effectively combines the advantages of diffusion models, regional convolutional neural networks (RCNNs), and transformers to handle dense objects with multi-class and multi-scale. Diffusion models ([24]) generate bounding box proposals from Gaussian noise, free from using pre-defined anchors; RCNNs efficiently crops dense instance maps into dense regions which has very high occupancy rate w.r.t. the bounding boxes (see Figure.2. (a)); Cross-attention mechanisms with dynamic queris [6, 29, 49, 14, 30] extract long-range contextual features to represent multi-scale objects robustly. Nevertheless, DiffRegFormer is not a simple composition of existing models, since we have observed that one composition can work in instance segmentation of sparse objects on datasets such as MSCOCO ([33]), but fail in the dense instance segmentation on the ROIs of kidney biopsies (please refer to ablation study 4.4.1). The reason behind this is that the proposals at the early training stage of diffusion models are almost pure Gaussian noises, which naturally sabotages the training due to accumulative errors. Therefore, we have made key innovations for dense instance segmentation:
  • 1.
    Regional features: Similar to RCNN, feature maps are cropped using generated proposals and converted to dynamic queries for efficient long-range dependency modeling, which is robust to vast variation in scales.
  • 2.
    Feature disentanglement: Instead of using shared feature maps, we redesign the model structure to generate feature maps for bounding box decoder and mask decoder seperately, which is crucial to stablize the training of mask decoder.
  • 3.
    Class-wise balanced sampling: In contrast to conventional sampling methods, we propose a novel sampling apporach. The key difference is to select class-wise balanced positive samples among the ground-truth boxes instead of the proposal ones.
In summary, the proposed model is a generic and extensible approach to large-scale datasets with little overhead, based on those key designs. To the best of our knowledge, DiffRegFormer is the first end-to-end framework that combines diffusion methodology with a transformer in RCNN style to efficiently process dense objects within ROIs for multi-scale and multi-class. In Figure.3, a flow chart of the instance segmentation is depicted.
For the lesion classifier, each anatomical structure has a dedicated classification head since class-wise lesions can be different depending on the lesion combinations in the dataset. In Figure.6, we input the crops of anatomical structures and resized them to the same size. Each head is trained only with crops of the specific structures and learns to predict multi-label lesions over each class. Our dataset currently only contains sclerotic glomeruli and atrophic tubuli, but the lesion classifier is extensible to expanded large-scale dataset at minimal cost, because adding, removing, or replacing heads cannot disrupte the other modules.
Fig. 2: regional segments vs. instance segments. (a) cropped instance maps from boxes tightly surrounding each object; (b) ground-truth bounding boxes and instance segments within one image; (c) separated instance mask per object.
The goal of our work is to propose a generic techniqical solution that has flexibility and scalability for automating lesion classification on dense anatomical structures in clinical scenarios, and work as the foundation for further quantification analyses. The main contributions of our work are the following:
  • 1.
    We propose the first end-to-end dense instance segmentation that effectively combines duffsion models, RCNNs, and transformers for multi-class, multi-scale objects within ROIs of kidney biopsies.
  • 2.
    We propose novel designs to address the accumulative errors caused by diffusion models at the early stage, and stablize the training process.
  • 3.
    We propose a novel class-wise balance sampling method to further improve the performance on detection and segmentation.
  • 4.
    Instead of the global queries used in transformers, we convert regional features into dynamic queries and model the long-range dependencies between regional features while avoiding performance deterioration on dense objects.
  • 5.
    We fairly compare DiffRegFormer with previous models that can process multi-class and multi-scale objects within ROIs in an end-to-end manner. Our model outperforms the previously published results on the evaluation of Jones’ silver-stained images (commonly used in nephropathology).
  • 6.
    We show that DiffRegFormer has potential of stain (domain)-agnostic detection for PAS-stained images without stain-specific fine-tuning.
  • 7.
    Our lession classifier can flexibly adapt to large-scale dataset with potential changes in lesion combinations at minimal overhead.
The remainder of the paper is organized as follows. In section 2, all relevant work is introduced and we show all improvements w.r.t. previous research. In section 3, we describe each component of our model in detail. In section 4, we demonstrate the extensive evaluation of the model including comparison experiments and ablation studies. In section 5, we describe the advantages and limitations of our pipeline and show possible follow-up research for the future.
Fig. 3: Our DiffRegFormer is a one-stage method. Instead of using anticipated anchors, we impose Gaussian noise on ground-truth boxes and generate a fixed-sized set of random bounding boxes. With feature maps extracted from the encoder, the bbox-decoder iteratively learns to denoise and predicts class-wise candidate boxes. Due to many unevenly distributed candidates, we propose a sampling module to discard negative samples and keep a proportion of balanced positive samples for fast convergence. Finally, we make final instance masks according to the given selected positive samples in the mask-decoder. For illustration simplicity, we only choose one object per class.
2 Related Work
Semantic Segmentation with Post-Processing: In the context of lesion classification of dense structures for renal biopsies, semantic segmentation with post-processing has drawn substantial attention. Previous studies ([23, 2, 40, 15]) treat anatomy-wise lesions (e.g. atrophic tubuli) as semantic classes alongside basic anatomical structures (e.g. arteries, glomeruli). An auxiliary border class is often integrated during the training phase to facilitate spliting semantic masks into distinct instances. Although this approach has demonstrated notable performance, it faces challenges in scalability and adaptability to large-scale datasets with potential changes in lesion composition: (1) Reliance on ad-hoc procedures (e.g. thresholding and morphological operations ([19])) for border generation reduces generalizability. (2) Due to Inflexibility to dataset change, any modification to lesion types necessitates model redesign and retraining. (3) Computational cost scales with lesion classes in dataset. (4) The difficulty in modeling long-range correlations between objects of varying scales due to spliting ROIs into tiles. In contrast, our end-to-end framework overcomes these limitations by directly modeling dense objects within ROIs using attention mechanisms.
Diffusion Model: Diffusion models, a class of deep generative models ([24, 36]), learn to approximate complex distributions through iterative denoising. Despite success in image generation ([25]), natural language processing ([31]), and audio processing ([37]), their application to dense instance segmentation remains limited. Initial attempts ([8, 20]) introduced diffusion models to segmentation tasks, but the adaptation to dense instance segmentation remains challenging. Our ablation studies suggest two primary challenges: (1) The potential dominance of the diffusion modules over feature learning in shared feature maps destabilize the training of instance prediction modules. (2) Accumulative errors from the diffusion module’s early training stages impair the accuracy of dense object predictions. Our approach introduces innovative strategies to overcome these challenges, enabling the first successful implementation of a diffusion model for dense instance segmentation.
Transformer: Transformer-based approaches introduce an anchor-free paradigm in end-to-end manner for instance segmentation, leveraging their capability to capture long-range dependencies and aggregate contextual information. In general, the dependency extraction process is called attention mechanism while each context feature is denoted as a query for object representation. Early transformer-based methods ([5, 12, 41]) rely on global queries to model all objects within a dataset. These approaches are incompatible with dense object scenarios due to the inevitable expansion of the global query set with dataset size. Recent works ([29, 49, 14, 30]) advance towards dynamic queries, allowing for temporary modeling of objects within a mini-batch. Thus it reduces complexity in processing dense objects. However, their instance segment representation remains inefficient for dense instance segmentation [^4] We conducted experiments on our dataset with both official codes. They required more than twice the GPU memory than ours on a batch of size 2 with 500 queries, which is not infeasible for applications with dense objects. due to the expensive representation of instance segments (Figure.2.(c)) with low occupation rate. Our DiffRegFormer addresses this by integrating region proposals with transformers for improved representation efficiency.
RCNN: The Regional Convolutional Neural Network (RCNN) framework ([18, 38]) effectively generates candidate regions (i.e. proposals) for dense objects. For instance segmentation, RCNN-based methods used mask modules to predict binary instance masks within proposals. Early variants ([21, 26, 27, 28]) relied on class-agnostic proposals followed by class-specific mask prediction, constrained by the limitations of predefined anchors. More recent approaches ([17]) have explored the mapping from proposals to global queries, employing cross-attention mechanisms for instance mask prediction. However, this paradigm faces scalability issues in dense object segmentation due to the computational overhead associated with global queries (refer to QueryInst in Table.1). Our method adopts dynamic queries, balancing computational efficiency and scalability for dense instance segmentation.
In summary, our work combines diffusion modeling and attention mechanisms into a RCNN-style framework, which addresses challenges of scalability, adaptability, and efficiency inherent in existing approaches for dense instance segmentation in renal biopsies.
3 Methods
3.1 Preliminaries
Diffusion model. Diffusion approaches ([43, 39]) is a family of deep generative models, inspired by the principles of non-equilibrium thermodynamics ([42, 25]). Their operation is conceptualized as $T$-step sequential process, iteratively transiting from an initial state $Z_{0}$ to a final state of pure noise $Z_{T}$. Diffusion forward involves the progressive addition of Gaussian noise at each transition. Due to the special properties of the Gaussian distribution ([24]), it is possible to directly sample any intermediate noisy at state $Z_{t}$:
$q(Z_{t}|Z_{0})=\mathcal{N}(Z_{t}|\sqrt{\bar{\alpha}_{t}}Z_{0},(1-\bar{\alpha}_{t}\mathbf{I}))$ (1)
where $Z_{0}$ is the original data, $Z_{t}\ (t\leq T)$ is the latent state at step $t$, and $\bar{\alpha}_{t}$ is the cumulative noise variance schedule ([8]). Notably, the diffusion forward process introduces no trainable parameters. The objective of a diffusion model is to learn the reverse process. A neural network, $f_{\theta}(Z_{t},t)$, approximates the recovery of the original distribution $Z_{0}$ from a noisy state $Z_{t}$. Training optimizes this network by minimizing a loss function to denoise at arbitrary intermediate state $t$:
$\mathcal{L}=\frac{1}{2}\|f_{\theta}(Z_{t},t)-Z_{0}\|^{2}$ (2)
where $Z_{0}$ represents the ground-truth and $Z_{t}$ is the noisy input. During inference, the model $f_{\theta}$ reconstructs from pure noise $Z_{T}$ to the predicted data $Z_{0}$ through a series of reverse updates ([43]) with a step-size $s$: $Z_{T}\rightarrow Z_{T-s}\rightarrow\cdots\rightarrow Z_{0}$. A detailed derivation is provided in ([43]).
3.2 Dense Instance Segmentation Model
Due to the iterative proposal generation, DiffRegFormer necessitates multiple model executions at inference, posing a computational burden if the entire model runs on the raw image at each step during training. To address this issue, we decouple the image encoder from the remaining modules. This strategy enables the encoder to extract multi-scale feature maps from the raw input image only once. Then subsequent modules operate for iterative refinement of box and mask predictions from an initial state of Gaussian noise conditioned on these deep features. Moreover, to facilitate the mask-decoder training, we introduce a novel sampling module that randomly selects positive samples from each class, ensuring a balanced distribution of class-wise regional features. This, thereby, facilitates the learning of object representations with lower frequencies in the dataset.
In summary, the DiffRegFormer comprises four components:
  • 1.
    Encoder: Processes the raw input image once to extract multi-scale feature maps for bbox-decoder and mask-decoder seperately.
  • 2.
    Bbox-decoder: Utilizes the extracted feature maps to iteratively refine bounding box predictions.
  • 3.
    Sampling Module: Stablizes the training by randomly selecting class-wise balanced positive samples for the mask-decoder .
  • 4.
    Mask-decoder: Iteratively refines mask predictions based on the sampled regional features.
3.2.1 Encoder
The image encoder extracts high-level feature maps from the raw image for subsequent decoders. It utilizes a ResNet pre-trained on ImageNet ([22]) as its backbone, followed by a Feature Pyramid Network (FPN) ([32]) to generate multi-scale feature maps. After FPN, a stack of $3\times 3$ convolutional operations is applied to produce two separate sets of feature maps, thereby decoupling the bbox-decoder and mask-decoder.
3.2.2 Bbox decoder
The bbox-decoder builds upon DiffusionDet ([8]). As shown in Figure.4(b), it firstly takes a set of random boxes and multi-scaled feature maps from the encoder as inputs. Dynamic queries (Figure.4(a)) then are initialized from a regional pooling operator (RoIAlign ([21])) and a feed-forward network (FFN) ([46]). These queries undergo iterative refinement through multiple stages. Each stage takes proposal boxes and dynamic queries from the previous stage, generating refined dynamic queries, box predictions, and class predictions for the next stage. Figure.4(c) illustrates one refinement stage in details. Firstly new regional features (cropped by RoIAlign) interact with dynamic queries (after a self-attention module) via dynamic convolution ([11]), emphasizing regions likely containing objects while suppressing others. Hereafter, the enhanced regional features are processed by FFNs to produce refined boxes, box-wise classifications, and refined queries respectively.
3.2.3 Sampling
Conventional RCNN models use intersection over union (IoU) to classify a proposal box as positive if its IoU with a ground-truth box is $\geq 0.5$ ([21]). This approach, however, is ill-suited for diffusion methods, particularly in early training stages when diffusion methods produce predominantly noisy. This can hinder convergence or even cause failure due to error accumulation. Further, conventional IoU-based methods are susceptible to class imbalances, biasing the learning process towards frequently occurring objects, thereby neglecting rare instances. We propose a novel sampling method to address these issues. We firstly replace proposal boxes with ground-truth boxes, ensuring a sufficient supply of positive samples. This is justified because the bbox-decoder’s objective is to predict boxes that closely approximate ground truth. We then divide ground-truth bounding boxes into groups based on their classes. Within each group, we randomly select up to $n$ boxes as positive samples without repetition. This sampling strategy offers three distinct advantages:
  1. 1.
    Rare Instance Learning: The mask-decoder can effectively learn representations of infrequently occurring instances whenever they appear in the input, mitigating the risk of being overwhelmed by more common classes.
  2. 2.
    Randomize Frequent Instance: Learning on a randomized subset of frequent instances prevents model overfitting.
  3. 3.
    Training Stability: Employing selected ground-truth boxes as positive samples promotes stability and accelerates mask training.
3.2.4 Mask decoder
The mask decoder refines mask predictions iteratively, utilizing regional features cropped from multi-scale feature maps based on boxes selected by the sampling module. Instead of using global queries ([12, 13]), we compute cross-attention between dynamic queries and positive regional features. This highlights effective queries and filters others, enhancing regional features by focusing on pixels within positive regions, thereby facilitating accurate per-pixel instance score prediction (Figure 5).
Fig. 4: The bbox-decoder takes as input multi-scale feature maps and a set of proposal boxes and outputs the prediction of classes and boxes iteratively. (a) The orange rectangle in (b) generates the initial dynamic queries via a RoIAlign pooling operator and a feed-forward network (FFN). (b) The module consists of a dynamic queries initialization head and multiple boxes refinement heads. (c) One box refinement module (one blue rectangle in (b)) inputs the dynamic queries and proposal boxes from the previous stage and generates predictions and refined dynamic queries for the next stage.
Fig. 5: The mask-decoder takes multi-scale feature maps and a set of positive proposal boxes to predict instance masks. The dynamic queries interact with regional features (i.e. cross-attention) and only highlight pixels residing in proposal boxes. Final instance masks are generated from the enhanced regional feature maps.
Fig. 6: The lesion classifier consists of 2 prediction heads for glomeruli and tubuli respectively. It takes crops of anatomical structures and outputs the probability of class-wise lesions.
Fig. 7: An illustration for a diffusion model at different stages. For simplicity, there are only 5 states with skip step as 1. State 4 is pure Gaussian noise while state 0 is the ground truth. (a) In one instance of training process, the diffusion model only randomly selects several denoise paths to train the box decoder; (b) In one instance of inference process, the diffusion model can alternately pick between denoise and corrupt path to generate multiple predictions, and ensemble them together for the final refined results.
3.3 Lesion classifier
The lesion classifier comprises multiple independent lesion prediction heads. Each head specializes in a specific anatomical structure (e.g., glomeruli) and processes cropped images of the specified structure to predict probabilities for multi-label lesions. Figure.6 illustrates this design. The architecture of a single head consists of three convolutional blocks followed by a max-pooling operator. Each block contains a $3\times 3$ convolutional, batch normalization, and ReLU activation. We currently implement two prediction heads: one for glomeruli (predicting sclerosis) and one for tubuli (predicting atrophy).
Importantly, the modular design of the instance segmentation component allows lesion prediction heads to operate as independent plug-ins. This modularity is crucial for seamlessly adapting the model to the expanding datasets, particularly in clinical scenarios. It allows for the straightforward integration or removal of prediction heads with minimal impact on the overall model architecture. For example, if new annotations for arteries become available, a dedicated prediction head can be trained to classify arterial lesions and integrated into the model without modifying existing components. Likewise, this design allows for the extension of the model’s diagnostic scope to include additional lesions within a given structure (e.g., tubuli) through the simple replacement of an existing prediction head with a newly trained one on expanded data. This design promotes the model’s versatility and long-term adaptability within clinical scenarios where the scope of identifiable lesions may increase.
3.4 Implementation Details
The entire model adopts distinct strategies for training and inference phases. During training, the DiffRegFormer learns to adeptly reconstruct ground truth boxes at state 00 from an arbitrary noisy state $t$ (Figure.7.(a)). Additionally, the sampling module provides class-balanced ground-truth boxes to facilitate the mask decoder’s training. Lastly, The lesion classifier operates on crops generated from ground-truth instance masks, with each prediction head specializing in classifying lesions specific to the corresponding anatomical structure within these crops. Conversely, during inference, the DiffRegFormer assembles multiple predictions to refine final results (Figure 7(b)). Similarly, The sampling module is deactivated, and all proposals are directly forwarded to the mask decoder for instance segmentation prediction. Likewise, the lesion classifier utilizes instance masks predicted by DiffRegFormer to crop structures. Lastly, each prediction head predicts probabilities for class-specific lesions within these cropped regions.
3.4.1 Training
We first construct noisy boxes from the ground truth. The DiffRegFormer is then trained to recover bounding boxes and iteratively generate instance masks from an arbitrary state $t$. The algorithm is detailed in Figure.3.2.4.
Diffusion forward. Ground-truth boxes are first concatenated with additional random boxes to aligned to a fixed length across images, accounting for the variability in instance counts. Gaussian noise is then introduced the noise scale governed by $\bar{\alpha}$ as per Eq.(1). Following ([36]), the scale varies with state $t$ using a monotonically decreasing cosine schedule. Notably, for optimal performance, the signal-to-noise ratio requires a relatively high signal scaling value compared to image generation tasks ([9, 24, 16, 10, 8]).
Training Loss. The box decoder takes noisy boxes as input and predicts category classifications and box coordinates. Given the one-to-many relationship between ground-truth and predicted boxes, a set-loss function is firstly employed to match predictions to ground-truths based on the lowest cost ([5, 45, 50]). Furthermore, the mask decoder takes sampled positive boxes as input and predicts instance masks. Due to one-to-one mapping, we then use a Binary Cross Entropy loss function. Finally, since multiple lesions can exist within one crop, a Binary Cross Entropy loss function is also appled to multi-label prediction within each head.
Sampling. Our implementation uses $3$ classes and a maximum of $N=500$ objects per image. we set up to $n=\frac{500}{3}\simeq 133$ instances per class. More specifically, positive ground-truth boxes are sampled without repetition, not exceeding the class-wise maximum. Moreover, we disregard absent classes within the input. For instance, if there are 2 glomeruli, 0 vessels, and 200 tubuli ground-truth boxes within an ROI, the sampling module then would select 2 glomeruli, 0 vessels, and 133 random tubuli as positive samples, without duplication.
3.4.2 Inference
During inference, the DiffRegFormer starts from a standard Gaussian distribution corresponding to the final state $T$ as defined in Eq.(1). Then a progressive denoising operation reverses the predictions to the initial state 00. Algorithm (as in Figure.3.4.2) outlines this process.
Inference steps. Each inference iteration involves two primary operations. Firstly, the box decoder processes noisy boxes generated from the previous state $\mathrm{t_{now}}$ and predicts both categories and box coordinates. Subsequently, Denoising Diffusion Implicit Models (DDIM) ([35, 43]) introduce noise into the previously predicted boxes. This step generates new noisy boxes for subsequent state $\mathrm{t_{next}}$, facilitating a progressive denoising process across different states. Notably, each iteration performs a single denoising operation, transitioning from state $\mathrm{t_{now}}$ to $\mathrm{t_{next}}$. This characteristic allows for the assembly of intermediate predictions from multiple denoising steps (i.e. $t_{i}\to 0$ where $i\in[1,T]$) refine the final outcome. However, to improve efficiency, DiffRegFormer performs inference only once, directly transitioning from pure Gaussian noise at state $T$ to state 00, to obtain the final predictions.
Box replacement. Predicted boxes can be categorized as either positive (scores above a threshold, containing objects of interest) or negative (scores below the threshold, arbitrarily located). Direct applying negative boxes into DDIM would significantly degrade the quality of newly generated noisy boxes. This issue stems from the fact that negative boxes, originating from corrupted boxes during the training phase, deviate significantly from the Gaussian distribution. To ensure consistency between training and inference phases, we substitute negative boxes with random boxes sampled from a Gaussian distribution.
Table 1: Overall evaluation result
Bounding Boxes Instances
AP AP50 AP75 APS APM APL AP AP50 AP75 APS APM APL
one-stage QueryInst_300 37.8 59.3 47.4 13.5 21.5 47.4 38.9 59.4 43.5 8.8 19.0 44.1
QueryInst_500 30.1 31.5 22.8 0.9 8.7 24.1 32.5 31.9 22.7 0.4 7.8 27.7
two-stage Mask-RCNN 49.0 68.6 56.2 18.0 24.4 59.2 45.5 69.0 53.5 16.4 21.5 54.4
Cascade Mask-RCNN 49.9 67.3 58.4 7.6 24.1 61.4 45.0 67.8 54.5 6.5 21.6 54.5
our model DiffRegFormer 52.1 71.1 57.7 15.4 27.4 61.4 46.8 71.6 52.8 14.1 23.6 54.2
  • 1.
    All models use ResNet-50 as a backbone network, and trained 40000 iterations on one NVIDIA GeForce RTX 3090 GPU with batch size 2. APS, APM, and APL correspond to AP on small, medium, and large-sized objects respectively.
Table 2: Per class evaluation result
Bounding Boxes - Glomeruli Instances - Glomueruli
AP AP50 AP75 APS APM APL AP AP50 AP75 APS APM APL
one-stage QueryInst_300 62.6 74.5 71.6 - - 63.7 56.1 74.5 68.1 - - 57.1
QueryInst_500 36.7 47.5 44.0 - - 37.6 35.9 47.6 43.6 - - 37.3
two-stage Mask-RCNN 69.4 85.8 78.6 - - 70.4 67.0 85.8 78.8 - - 68.1
Cascade Mask-RCNN 70.7 84.0 81.7 - - 72.1 65.1 84.0 79.8 - - 66.4
our model DiffRegFormer 80.9 94.2 89.8 - - 80.9 74.3 94.2 80.9 - - 74.3
Bounding Boxes - Arteries Instances - Arteries
AP AP50 AP75 APS APM APL AP AP50 AP75 APS APM APL
one-stage QueryInst_300 18.8 37.0 19.1 5.2 20.2 20.6 17.5 36.3 14.8 4.5 17.9 19.4
QueryInst_500 4.5 10.0 2.9 0.5 5.6 4.5 4.7 10.5 3.8 0.1 5.3 5.3
two-stage Mask-RCNN 22.5 40.3 23.5 20.2 19.4 28.9 20.1 41.4 18.0 20.2 17.4 25.4
Cascade Mask-RCNN 23.0 38.0 26.9 - 17.8 33.3 20.1 39.4 21.9 - 16.9 27.5
our model DiffRegFormer 25.7 46.9 27.8 13.5 23.9 31.6 23.4 48.8 20.2 20.2 20.4 30.1
Bounding Boxes - Tubuli Instances - Tubuli
AP AP50 AP75 APS APM APL AP AP50 AP75 APS APM APL
one-stage QueryInst_300 45.8 66.4 51.4 21.8 44.3 57.9 41.2 67.3 47.6 13.1 39.0 55.8
QueryInst_500 21.2 37.1 21.7 1.3 20.3 30.1 20.6 37.5 20.7 0.7 18.1 40.5
two-stage Mask-RCNN 57.5 79.6 66.4 15.8 53.8 78.2 51.4 79.9 62.3 12.7 47.3 69.8
Cascade Mask-RCNN 58.1 79.9 66.5 15.1 54.4 78.8 51.8 80.0 61.7 13.0 47.9 69.6
our model DiffRegFormer 58.5 80.2 67.9 19.9 51.0 75.4 55.8 80.6 62.3 16.6 45.6 64.5
  • 1.
    All models use ResNet-50 as a backbone network, and trained 40000 iterations on one NVIDIA GeForce RTX 3090 GPU with batch size 2. APS, APM, and APL correspond to AP on small, medium, and large-sized objects respectively.
4 Results
In this section, we start with an overview of the kidney biopsy dataset used in this study. We then proceed to draw a fair comparison between DiffRegFormer against established end-to-end instance segmentation models specifically designed to handle dense, multi-class, and multi-scale objects at the ROI level. To ensure consistency, all benchmark methods are reproduced utilizing the mmdetection package ([7]). Further, we evaluate the performance of our lesion classifier on cropped images of anatomical structures, derived from the predicted instance masks by DiffRegFormer. Furthermore, we conduct a complexity analysis of DiffRegFormer to identify an optimal balance between its performance and computational efficiency. Finally, we present extensive ablation studies on DiffRegFormer to evaluate the impact of various training strategy combinations.
4.1 Datasets
Biopsy specimens were prepared in accordance with the Pathology Laboratory Protocol for kidney biopsies. Tissues were collected via a standard procedure of operations: core needle biopsy, formalin-fixed, rapidly processed by a routine tissue processor, embedded in paraffin, serially sectioned at $3-4\mu m$, mounted onto adhesive slides, and stained with Jones silver staining. The dataset comprises $148$ patient biopsy samples (both transplant and native kidneys) collected from the multicenter’s archive of the Departments of Pathology at LUMC, AUMC, UMCU in the Netherlands or the archive of the Department of Nephrology at Leuven in Belgium. Prior to analysis, all biopsies were anonymized by a pathology staff member. Annotation was performed employing ASAP version $1.9$ for Windows [^5] https://computationalpathologygroup.github.io/ASAP/ and our custom software Slidescape [^6] https://github.com/amspath/slidescape. The $148$ Jones-stained WSIs were acquired from scanners at a resolution of $\mathrm{PPM}=0.25$ in BIG-TIFF format [^7] https://www.awaresystems.be/imaging/tiff/bigtiff.html. Openslide [^8] https://openslide.org/ was used to extract $303$ ROIs at level 00 of the BIG-TIFF images. Table.3 outlines the dataset composition, and Figure.8 illustrates the distribution of ROI short sizes. Additionally, $115$ PAS-stained WSIs from LUMC were utilized solely for testing of domain transferability. Notably, ROIs from the same WSIs were exclusively allocated to either training or validation. To train the lesion classifier, $21,889$ patches of individual objects were extracted from $303$ ROIs, including $916$ glomeruli patches with binary lesion labels ($382$ with sclerotic and $534$ healthy) and $20,973$ tubuli patches with binary labels ($10,647$ with atrophy and $10,326$ healthy), as detailed in Figure.6.
Table 3: Instance Data Set Configuration.
Source Scanner WSI ROIs Train Eval
LUMC Philips UFS 65 142 130 12
AUMC Philips UFS 42 80 74 6
UMCU Hamamatsu XR 25 49 45 4
Leuven Philips UFS 16 32 - 32
Fig. 8: The histogram of the shortest edges of ROIs extracted from WSIs. In our dataset, the shortest length is 387 and the largest one is 4565.
Fig. 9: A visualization for all methods with Jones-stained kidney biopsies. Our model can detect objects of various sizes and shapes. Over each region, decent instance masks can be generated although there are some overlays between different instances. Mask-RCNN-based models cannot locate one large object within one box because they do not adopt attention mechanisms to capture long-range dependencies. QueryInst with global queries cannot process dense objects since they fail to detect every object within one biopsy.
Fig. 10: A visualization for our model on the PAS-stained kidney biopsies with increasing degrees of fibrosis, inflammation (left to right), and variation in stain intensity.
Fig. 11: Complexity Analysis on Jones-stained datasets.
(a) shared feature + proposals + biased sampling on MSCOCO’s image
(b) shared feature + ground-truth + biased sampling on a ROI
(c) shared feature + ground-truth + unbiased sampling on a ROI
(d) separate feature + proposals + unbiased sampling on a ROI
(e) separate feature + ground-truth + biased sampling on a ROI
Fig. 12: Ablation study on combinations of different strategies.
Table 4: Evaluation on combinations of different strategies
feature boxes sample glo-det art-det tub-det all-det glo-ins art-ins tub-ins all-ins
75.8 24.9 56.4 52.4 - - 33.6 -
76.1 26.2 57.6 53.3 7.3 4.8 35.2 15.8
8.5 4.3 13.5 8.8 6.8 4.1 12.5 7.8
12.6 6.8 56.7 25.4 10.4 5.3 45.8 20.5
80.9 25.7 58.5 55.1 74.3 23.4 55.8 51.2
  • 1.
    glo: glomeruli. art: arteries. tub: tubuli. all: overall. det denotes detection. ins means instance. ”-” means not working.
  • 2.
    feature: if use seperate feature maps between bbox-decoder and mask-decoder. boxes: if use ground-truth boxes in the sampling process. sample: if maintain class balance in the sampling process.
  • 3.
    The symbol ✓ denotes taking one strategy while ✗ means taking the converse strategy.
Table 5: Evaluation of detection performances on different types of queries
# 200 # 300 # 400 # 500 # 600
dynamic 33.6 39.4 45.8 52.1 51.7
global 23.8 38.9 37.4 34.6 31.6
Table 6: Lesion data set configuration and classification Performance
tubuli glomeruli
atrophy normal sclerosis normal
train 10647 10326 382 534
eval 568 906 22 51
per_precision 85.1 93.2
per_recall 40.5 88.7
all_precision 89.2
all_recall 64.6
Fig. 13: A visualization of ablation study on query types.
Fig. 14: A visualization of DiffRegFormer predicts dense objects on one Whole Slide Image (WSI).
Fig. 15: A visualization of lesion classifier predicts the probability of class-wise lesion on ROIs.
4.2 Experiments
In our experimental analysis, we conducted a fair comparison of our proposed DiffRegFormer against well-established end-to-end instance segmentation models: Mask R-CNN ([21]), Cascade Mask R-CNN ([4]), and QueryInst ([17]), specifically designed to simultaneously handle dense, multi-class, multi-scale objects at the ROI level. All the models under comparison employed a ResNet-$50$ backbone pre-trained on the ImageNet ([22]) and were trained on $249$ ROIs over $40000$ iterations. We utilized the mmdetection data augmentation pipeline: each ROI was first subjected to RandomFlip with a $0.5$ probability, followed by random selection between RandomResize or RandomCrop also with a probability $0.5$. RandomResize adjusted the ROI to a size with the shortest edge between $480$ and $1333$ pixels, while RandomCrop extracted patches with the shortest edge between $384$ and $600$ pixels. Finally, pixel values were normalized to $(-1,1)$. As mmdetection supports training with variable image sizes, all models were allowed to train on objects at multiple scales.
For feature extraction, we used $256$ channels for queries. During both training and inference, DiffRegFormer employed 500 dynamic queries, while QueryInst used $300$ and $500$ global queries respectively. Cascade Mask-RCNN, QueryInst, and our method utilized $6$ cascading stages. Table.1 presents performance comparisons for object detection w.r.t bounding boxes and instance segmentation w.r.t masks respectively. For all experiments, DiffRegFormer did not utilize refinement during inference (i.e., iteration set to 1). Notably, Mask-RCNN ([21]) and cascade Mask-RCNN ([4]) are two-stage methods employing RPN networks instead of dynamic queries. QueryInst is a one-stage method with global queries. In contrast, DiffRegFormer is categorized as a one-stage method with dynamic queries.
Detection. DiffRegFormer achieves a detection Average Precision (AP) of 52.1% with a ResNet-50 backbone, surpassing competitors such as QueryInst, Mask R-CNN, and Cascade Mask R-CNN by notable margins. Intriguingly, iterative refinement appears to hinder detection performance for small objects while benefiting the detection of medium to large objects. This suggests the need for additional operations to mitigate information loss for small objects during the refinement process. Furthermore, the use of global queries appears less effective for handling dense objects, particularly when the number of queries significantly exceeds the number of feature channels ($300$ or $500$ compared to $256$), leading to a marked decline in detection performance.
Instance segmentation. DiffRegFormer achieves a 46.8% AP for instance segmentation, outperforming its counterparts. Likewise, the application of iterative refinement to bounding boxes is found to significantly impact the instance segmentation outcomes. This effect can be attributed to the binary segmentation process within the boxes, where the quality of the instance masks is inherently linked to the recall and precision rates of the bounding boxes.
Table.2 presents the performance per object class. Our model demonstrates an overall superior ability to detect all anatomical structures compared to other methods. Additionally, it excels in the instance segmentation of glomeruli and arteries by notable margins and shows a slight advantage in tubuli segmentation. These findings indicate our model’s ability to process dense, multi-class objects with diverse scales at the ROI level. However, we observe a performance bottleneck for all models in the processing of arteries, which poses a limitation for clinical application due to the significant size variation within the dataset. For instance, the largest arteries are hundreds of times larger than the smallest. This fact suggests the need for more sophisticated attention mechanisms to capture global contextual information for such extreme scale variations.
Figure.9 demonstrates DiffRegFormer’s capability to effectively detect objects at various scales. It accurately generates instance masks within bounding boxes, demonstrating precise localization of larger arteries (see the second row of the figure). In contrast, Cascade Mask R-CNN has difficulty in locating larger objects within a single bounding box, due to the lack of an attention mechanism for capturing long-range dependencies. Furthermore, Mask-RCNN tends to generate multiple detections for a single object(as seen in the third row), resulting from the absence of cascade refinement. These observations suggest that iterative bounding box refinement during training significantly reduces multiple predictions for a single object, while the attention mechanism with dynamic queries processes large-scale objects. Nevertheless, QueryInst, with its reliance on global queries, may struggle to detect all instances effectively in datasets with significant size and shape variations. QueryInst-$300$ tends to miss many structures, while QueryInst-$500$ can even degenerate into generating multiple false predictions for a single object. Furthermore, RCNN-based models share a limitation: a single pixel may be allocated to multiple instance masks due to pixel-wise predictions within each independent ROI. This inherently prevents enforcing the spatial exclusivity of pixels to instances. Additionally, Figure.10 demonstrates DiffRegFormer’s potential for domain transfer, as it generates reasonable results on PAS-stained images despite being trained exclusively on Jones’ silver-stained images. Finally, Figure.14 highlights that DiffRegFormer, when combined with techniques such as sliding window and stitching, can be effectively applied to WSIs, further demonstrating its effectiveness.
Lesion classification. The data augmentation pipeline for training the lesion classifier is relatively straightforward. We set the batch size as $96$ for the input crops, where $32$ crops for each class. Each crop first undergoes RandomFlip with a probability of $0.5$, followed by resizing to $256\times 256$. Finally, pixel values are normalized to the range of $(-1,1)$. As shown in Table.6, the classifier achieves an overall precision of $89.2\%$ precision and a recall of $64.6\%$. More specifically, per-class precision for tubuli atrophy is $85.1\%$, and $93.2\%$ for glomeruli sclerosis. In addition, per-class recall is $40.5\%$ for tubuli atrophy and $88.7\%$ for glomeruli sclerosis. Since the lesion classifier operates on crops of anatomical structure generated by DiffRegFormer during inference, its recall rate is strongly influenced by the performance of our dense instance segmentation model. In contrast, the high precision indicates that lesion classification on correctly cropped structures is highly accurate. In summary, enhancing the performance of DiffRegFormer is expected to further improve the lesion classifier’s recall rate. Figure.15 visualizes the mapping of lesion predictions onto the instance segmentation results.
4.3 Complexity Analysis
We start with analyzing the class imbalance within our dataset. Of the total structures ($21,889$ objects), 88% are tubuli, followed by 7% arteries (1679 objects), and 5% glomeruli (1226 objects). This distribution is visualized in Figure.11(a), highlighting the necessity of sampling to ensure the inclusion of rare instances, such as arteries and glomeruli, among the dominant tubuli. A violin plot, shown in Figure.11(b), details the distribution of structure counts at the ROI level. On average, ROIs contain $95$ objects, with a maximum of $434$. To accommodate all objects, DiffRegFormer is configured to handle a maximum number exceeding $434$ per image. The correlation between the increase in computational cost and performance improvement, depicted in Figure 11(c). As expected, we observe a turning point at $500$ dynamic queries. This configuration yields a favorable trade-off between complexity and efficiency, achieving 52.1% AP with $20.2$GB GPU memory usage. Specifically, ROIs contain an average of $5$ glomeruli, $7$ arteries, and $84$ tubuli with corresponding maximums of $74$, $84$, and $408$. Setting the maximum number per class to $133$ is another feasible option for positive sample generation, as our sampling approach ensures comprehensive inclusion of rare instances while maintaining rational proportion of the majority class. In conclusion, selecting a maximum of $500$ objects per image and $133$ per class achieves an advantageous balance between performance and computational efficiency.
4.4 Ablation Study
In this ablation study, we focus on the effects of various training strategy combinations and query types on model performance. Through analysis of these results, we can gain insights into the specific function of each strategy. This analysis highlights the significance of our improvements as key technical contributions for adapting an end-to-end instance segmentation model to datasets with dense objects.
4.4.1 Training strategies
Training an instance segmentation model for datasets with dense objects, such as kidney biopsies, presents distinct challenges compared to datasets with sparser object distributions like MSCOCO. Instance segmentation models typically employ three strategies: (1) shared features between the bbox-decoder and mask-decoder; (2) selection of positive samples from proposal bounding boxes; and (3) class-imbalanced sampling. However, for dense object segmentation, we propose three alternative strategies: (a) separate features for the bbox and mask decoders; (b) positive sample selection directly from ground truth boxes; and (c) class-balanced sampling. While the conventional combination of shared features, proposal boxes, and biased sampling is effective for standard instance segmentation (Figure.12a), only the combination of separate features, ground-truth boxes, and unbiased sampling proves successful for dense instance segmentation. Table.4 presents an evaluation of these combinations, with conventional strategies marked with ✗ and our proposed strategies with ✓. Notably, only five out of eight possible combinations can train DiffRegFormer.
Figure.12b demonstrates that the combination of shared features, ground-truth boxes, and biased sampling enables object detection across all classes. However, it fails to produce accurate instance masks within these bounding boxes. This suggests that the bbox decoder dominates the shared feature representation, hindering the mask decoder’s ability to extract pixel-wise information. Consequently, all pixels within detected boxes are classified as tubuli. Moreover, biased sampling impedes the mask decoder’s capacity to learn features for rare instances (glomeruli and arteries), resulting in no pixel classification within those boxes.
Additionally, Figure.12c, shared features combined with ground-truth boxes and unbiased sampling can detect and segment objects across all classes, albeit with the degraded mask quality. This degradation is likely due to the bbox decoder’s dominance within the shared feature representation.
In contrast, Figure.12d reveals that using separate features, proposal boxes, and unbiased sampling leads to inaccurate object detection. This is likely due to the accumulation of errors introduced by proposals during the early stages of diffusion model training.
Lastly, Figure.12e demonstrates that separate features, ground-truth boxes, and biased sampling result in correct detection and segmentation for only a small proportion of tubuli. This finding further highlights the challenges posed by biased sampling in learning features for rare instances.
In summary, shared feature maps lead to the bbox-decoder dominating feature representation learning, impairing the mask-decoder’s ability to extract pixel-wise information necessary for accurate binary instance masks. Biased sampling further hinders the mask-decoder to learn information for segmenting rare instances (glomeruli and arteries), causing it to be overwhelmed by the majority class (tubuli). Finally, utilizing proposals for sampling instead of ground truth boxes introduces cumulative errors that disrupt the training for both decoders, preventing to learn contextual information necessary for long-range correlations via attention mechanisms.
4.4.2 Queries
Global queries in transformer-based models face challenges for representing dense objects within large-scale datasets due to their static nature. Effectively accommodation to extensive variations in object size and shape necessitates a proportional increase in query count. Conversely, dynamic queries, generated on-the-fly for each object in a mini-batch, demonstrate superior adaptability to dataset complexity. Theoretically, assigning a larger number of global queries could achieve comparable performance to dynamic queries. However, our experiments demonstrate that increasing the number of global queries without a corresponding growth in feature channels does not yield improvements. Consequently, employing global queries for dense object processing in transformer-based models is significantly less cost-effective than utilizing dynamic queries.
Table.5 delineates the comparative analysis of varying query counts for both dynamic and global queries, with the number of feature channels fixed at $256$. We observe that increasing the number of dynamic queries leads to performance gains, while increasing the number of global queries results in performance degradation. Figure.13 further illustrates the inherent challenge of achieving a favorable trade-off between complexity and performance with global queries. Even with $500$ queries, the model with global queries fails to consistently detect structures across all classes, a limitation not attributable to class imbalance, in contrast to the dynamic query strategy. In conclusion, using dynamic query is a more efficient approach for datasets with dense objects, highlighting their superiority over global queries.
5 Conclusion and Future Work
This research introduces a novel framework to process large-scale datasets with potential changes in lesion combinations. Our model demonstrates superior performance in predicting lesions within dense objects, effectively handling multi-class and multi-scale challenges at the ROI level. We present the first dense instance segmentation module, DiffRegFormer, that seamlessly integrates a diffusion model with a transformer in RCNN style. This approach leverages the diffusion model to generate refined object boxes, upon which a regional transformer creates dynamic queries that are ultimately transformed into accurate instance masks within the boxes. Furthermore, our lesion classifier accurately predicts class-specific lesions for each cropped structure. The model demonstrates several advantages in handling dense object datasets: (1) its iterative noise-to-box approach eliminates reliance on prior knowledge about object size and shape, facilitating rapid object localization; (2) dynamic queries, generated on-the-fly for each object in a mini-batch, ensures scalability to large datasets; (3) the model’s modular design, with independent plug-in prediction heads, allows for replacement and promotes adaptability. This flexibility enables efficient reuse of trained components when scaling to large-scale datasets. Experimental results on Jones’ stained renal biopsies demonstrate that our model surpasses existing benchmarks.
While the current model may not yet meet the rigorous requirements for clinical applications, it establishes a robust framework for dense, multi-class, multi-scale object recognition at the ROI level, setting a solid foundation for future engineering improvements. Potential areas for future research include: (1) further enhancing the compatibility of diffusion models with transformers. (2) developing innovative regional feature extractors to generate more informative dynamic queries and mitigate the loss of information for small objects during iterative refinements. (3) integrating spatial constraints into regional mask inference to prevent erroneous assignment of single pixels to multiple instances.
6 Acknowledgements
This work is supported in part by funds from the Dutch Kidney Foundation (17OKG23), the Research Priority Area Human(e) AI at the University of Amsterdam, and the Chinese Scholarship Council (CSC).